OpenStudy (anonymous):

The length of a rectangle is four times its width. If the area of the rectangle is 400ft^2 , find its perimeter.

OpenStudy (anonymous):

106.8 is incorrect

OpenStudy (nurali):

If the area of the rectangle is 400ft, find its perimeter. L = 4W Equation: Area = LW 400 = (4W)W 100= W^2 W = 10 L = 4W = 40 Perimeter = 2(L+W) P = 2(40+10) P = 2(50) P = 100

OpenStudy (nurali):

if correct please click best response.